OpenStudy (anonymous):

logbase2(x) +4logbase2(y)-3logbase2(z) can someone rewrite as a single logarithm logbase2 A please

OpenStudy (jdoe0001):

using the logarithmic rule that \( log_b\alpha^\beta \implies \beta*log_b\alpha \), convert the 2nd and 3rd terms

OpenStudy (anonymous):

can you help me please i dont know how

OpenStudy (jdoe0001):

well, look at what you have for 2nd and 3rd terms

OpenStudy (anonymous):

4 and -3

OpenStudy (jdoe0001):

not just the coefficients, the term itself, literals and all

OpenStudy (anonymous):

4logbase2(y)-3logbase2(z) 4logbase2(y)-3logbase2(z)

OpenStudy (jdoe0001):

\( \beta*log_b\alpha \implies (4)log_2Y \ \&\ (-3)log_2Z \)

OpenStudy (anonymous):

is that the solution

OpenStudy (jdoe0001):

and the rules is \(\beta*log_b\alpha \implies log_b\alpha^\beta \)

OpenStudy (jdoe0001):

no, is not the solution, you need to convert them first, to get the the solution

OpenStudy (anonymous):

ok thanks

OpenStudy (jdoe0001):

http://www.adug.org.au/MathsCorner/LOGS1.GIF